OpenStudy (vivi2827):

Can you please help me with this equation 2/w=x/5 solve for w

Awolflover1 (awolflover1):

Multiply both sides by w

Awolflover1 (awolflover1):

\[2=\frac{ wx }{ 5 }\]

Awolflover1 (awolflover1):

now Multiply both sides by 5 10=wx

Awolflover1 (awolflover1):

then flip it \[wx=10\]

Awolflover1 (awolflover1):

Divide both sides by x

OpenStudy (mathmale):

Solve for w? 2/w=x/5 Consider the following approach:\[ \frac{ 2 }{ w }=\frac{ x }{ 5}\]

OpenStudy (mathmale):

Invert both sides, obtaining:\[\frac{ w }{ 2 }=\frac{ 5 }{ x }\]

OpenStudy (mathmale):

It's now easy to solve for w. What would your last step be?

jhonyy9 (jhonyy9):

so the best easy way to solve it for w cross multiplie and will get 10=wx divide both sides by x and will get 10/x = w hope helped